# On the power of graph neural networks and the role of the activation function

### v4

@@ 4. Collision with piecewise polynomial activations
where $\phi$ is a piecewise bivariate polynomial. By composition $\xi^{t+1}(T[k_{1},\ldots,k_{m}],v)$ is a piecewise multivariate polynomial of $k_{1},\cdots,k_{m}$. By induction, the degree of $\xi^{t+1}(T[k_{1},\ldots,k_{m}],v)$ depends only on $t$ and the degree of $\phi$, which is independent of $m$. To conclude, we will need the following:
Claim: Let $T[k_{1},\cdots,k_{m}]$ be two rooted trees given by Figure 1, and let $s$ be its source vertex. Then for any iteration $t$, the color refinement embedding $\mathsf{cr}^{t}(s)$ is invariant by permutation of the $k_{i}$’s.
Proof.
By induction of $t\geq 0$, we prove that for any integer $t\geq 0$, $\mathsf{cr}^{t}(s)$ and the multiset $\{\{\mathsf{cr}^{t}(x_{1}),\allowbreak\cdots,\mathsf{cr}^{t}(x_{m})\}\}$ are invariant by permutation of the $k_{i}$’s.
Base case: $(t=0)$ is obvious since $\mathsf{cr}^{0}(v)=1$ for any vertex $v\in T[k_{1},\cdots,k_{m}]$.
Induction step: Suppose that for some iteration $t$, $\mathsf{cr}^{t}(s)$ and $\{\{\mathsf{cr}^{t}{x_{1}},\cdots,\mathsf{cr}^{t}(x_{m})\}\}$ are invariant by permutation of the $k_{i}$’s. $\mathsf{cr}^{t+1}(s)$ being invariant by permutation of the $k_{i}$’s is a consequence of: $\mathsf{cr}^{t+1}(s)=(\mathsf{cr}^{t}(s),\{\{\mathsf{cr}^{t}(x_{1}),\cdots,\mathsf{cr}^{t}(x_{m})\}\})$
Finally, we know from [Gro21, Theorem VIII.1] that the color refinement algorithm refines any GNN at any iteration. Since the tuple obtained by color refinement for the vertex $s$ is invariant with respect to permutations of the $k_{i}$’s, $\xi^{t}(T[k_{1},\ldots,k_{m}],s)$ is also invariant with respect to permutations of the $k_{i}$’s. ∎
@@
Let $T>0$ be a positive integer, and for $0\leq t\leq T$, let $f_{t,m}(k_{1},\cdots,k_{m}):=\xi^{t}(T[k_{1},\ldots,k_{m}],s)$ be the value returned by a GNN with piecewise polynomial activation after $t$ iterations (note that the embeddings are one-dimensional because only one color is used). Due to Lemma 3, for any $t\in\{0,\cdots,T\}$, $((k_{1},\cdots,k_{m})\mapsto f_{t,m}(k_{1},\cdots,k_{m}))_{m\in\mathbb{N}}$ is a sequence of symmetric piecewise multivariate polynomials with bounded degrees (the degree of $f_{t}$ is upper-bounded by a constant that does not depend on $t$ nor $m$). Lemma 2 with $I=\{0,\cdots,T\}$ tells us that there exists $m\in\mathbb{N}$, and two vectors $k\in\mathbb{N}^{m}$ and $k^{\prime}\in\mathbb{N}^{m}$ whose coordinates are not equal up to permutations, such that for any $t\in\{0,\cdots,T\}$, $f_{t}(k_{1},\cdots,k_{m})=f_{t}(k^{\prime}_{1},\cdots,k^{\prime}_{m})$. ∎
@@ 6. Acknowledgements
The authors gratefully acknowledge support from Air Force Office of Scientific Research (AFOSR) grant FA95502010341 and National Science Foundation (NSF) grant CCF2006587. The authors are also very grateful to insightful comments from reviewers that helped to improve the paper significantly. In particular, we learnt of the reference [${}^{+}$22] from one of the reviewers.

### v5

@@ 4. Collision with piecewise polynomial activations
where $\phi$ is a piecewise bivariate polynomial (as a Neural Network $\mathbb{R}^{2}\rightarrow\mathbb{R}$ with piecewise polynomial activation). By composition $\xi^{t+1}(T[k_{1},\ldots,k_{m}],v)$ is a piecewise multivariate polynomial of $k_{1},\cdots,k_{m}$ such that for every integer $m$, $\mathsf{deg}(\xi^{t+1}(T[k_{1},\linebreak\ldots,k_{m}],v))\leq q_{t}\cdot\mathsf{deg}(\phi)$. Furthermore, since for every $m$ and every vertex $u$, $\xi^{t}(T[k_{1},\ldots,k_{m}],u)$ is supposed to be $q_{t}$ generated, then using the above update rule, Lemma 2 gives us that by composition, $\xi^{t+1}(T[k_{1},\ldots,k_{m}],v)$ is $\tilde{q_{t}}$ generated, where $\tilde{q_{t}}$ depends only on the degre of $\phi$ and $q_{t}$.
Setting $q_{t+1}:=\max(q_{t}\cdot\mathsf{deg}(\phi),\,\tilde{q_{t}})$ gives us the desired properties (for every vertex $v$ and every integer $m$, $\xi^{t+1}(T[k_{1},\cdots,k_{m}],v)$ has degree at most $q_{t+1}$ and is $q_{t+1}$ generated), and ends the induction on $t$. ∎
@@
Let $T>0$ be a positive integer, and for $0\leq t\leq T$, let $f_{t,m}(k_{1},\cdots,k_{m}):=\xi^{t}(T[k_{1},\ldots,k_{m}],s)$ be the value returned by a GNN with piecewise polynomial activation after $t$ iterations (note that the embeddings are one-dimensional because only one color is used). Using Lemma 3, there exists an integer $q$ such that the double sequence $(f_{t,m})_{t\in\{0,\cdots,T\},m\in\mathbb{N}}$ of piecewise multivariate polynomials has degree at most $q$ and such that every $f_{t,m}$ is $q$-generated. Lemma 1 with $I=\{0,\cdots,T\}$ tells us that there exists $m\in\mathbb{N}$, and two vectors $k\in\mathbb{N}^{m}$ and $k^{\prime}\in\mathbb{N}^{m}$ whose coordinates are not equal up to permutations, such that for any $t\in\{0,\cdots,T\}$, $f_{t,m}(k_{1},\cdots,k_{m})=f_{t,m}(k^{\prime}_{1},\cdots,k^{\prime}_{m})$. ∎
@@ 6. Acknowledgements
The authors gratefully acknowledge support from Air Force Office of Scientific Research (AFOSR) grant FA95502010341 and National Science Foundation (NSF) grant CCF2006587. The authors are also very grateful to insightful comments from reviewers that helped to improve the paper significantly. In particular, we learnt of the reference [ACI+22] from one of the reviewers. We also thank Eran Rosenbluth for pointing out an error in our first proof of Theorem 3.